A car with a mass of 2000 kg is moving around a banked curve with a constant speed of 20 m/s (about 44 MPH). The radius of curvature of the curve is 73 m. (a) What is the centripetal acceleration of the car? (b) What is the magnitude of the horizontal component of the normal force that would be required to produce this centripetal acceleration in the absence of any friction?
Answer #1

given

Mass of car m = 2000 kg
velovity v =20 m/s
Radius of curvature r = 73 m

a) centripetal acceleration, a = v^2/r

= 20^2/73

= 5.48 m/s^2

b) Horizontal component of normal force = m*a

= 2000*5.48

= 10958.9 N